# TabArena: A Living Benchmark for Machine Learning on Tabular Data

### v2

@@ D.5 Running TabArena Models in Practice
Appendix E Performance Results Per Dataset
Table E.1 presents the performance per dataset for all methods in TabArena-v0.1.
@@ Appendix E Performance Results Per Dataset
Table E.1: Performance Per Dataset.  We show the average predictive performance per dataset with the standard deviation over folds.  We show the performance for the default hyperparameter configuration (Default), for the model after tuning (Tuned), and for the ensemble after tuning (Tuned + Ens.).  We highlight the best-performing methods with significance on three levels:  (1) Green: The best performing method on average;  (2) Bold: Methods that are not significantly worse than the best method on average, based on a Wilcoxon Signed-Rank test for paired samples with Holm-Bonferroni correction and $\alpha=0.05$.  (3) Underlined: Methods that are not significantly worse than the best method in the same pipeline regime (Default, Tuned, or Tuned + Ens.), based on a Wilcoxon Signed-Rank test for paired samples with Holm-Bonferroni correction and $\alpha=0.05$. We exclude AutoGluon for significance tests in the Tuned + Ens. regime.
@@
RF 4889.0 $\pm$ 289.0 4641.0 $\pm$ 304.0 4630.0 $\pm$ 303.0
ExtraTrees 4845.0 $\pm$ 275.0 4607.0 $\pm$ 324.0 4610.0 $\pm$ 323.0
XGBoost 4672.0 $\pm$ 306.0 4523.0 $\pm$ 320.0 4520.0 $\pm$ 320.0
LightGBM 4610.0 $\pm$ 314.0 4525.0 $\pm$ 329.0 4512.0 $\pm$ 325.0
CatBoost 4535.0 $\pm$ 329.0 4518.0 $\pm$ 322.0 4519.0 $\pm$ 321.0
EBM 4549.0 $\pm$ 319.0 4500.0 $\pm$ 333.0 4499.0 $\pm$ 326.0
FastAIMLP 4720.0 $\pm$ 313.0 4634.0 $\pm$ 318.0 4624.0 $\pm$ 311.0
TorchMLP 4661.0 $\pm$ 343.0 4526.0 $\pm$ 329.0 4534.0 $\pm$ 333.0
RealMLP 4579.0 $\pm$ 313.0 4571.0 $\pm$ 324.0 4535.0 $\pm$ 323.0
TabM 4514.0 $\pm$ 323.0 4530.0 $\pm$ 323.0 4511.0 $\pm$ 326.0
MNCA 4605.0 $\pm$ 332.0 4614.0 $\pm$ 330.0 4589.0 $\pm$ 331.0
TabPFNv2 4695.0 $\pm$ 303.0 4650.0 $\pm$ 337.0 4568.0 $\pm$ 319.0
TabDPT 4508.0 $\pm$ 295.0 - -

### v3

@@ E.5 Running TabArena Models in Practice
E.6 Handling Foul Play and Dataset Contamination
A fundamental limitation of open-source benchmarking is that foul play and dataset contamination could compromise the leaderboard of TabArena. Model developers could overfit the hyperparameters of their model on the TabArena datasets, or use them for pretraining a foundation model.
Below, we provide a discussion on handling foul play and benchmarking foundation models with dataset contamination. In the future, we aim to explore various solutions to these challenges with TabArena to keep our leaderboard representative for practitioners.
Avoiding Foul Play. Foul play will inevitably affect TabArena. Thus, we, as maintainers, have considered future guards against foul play in four ways:
  1. 1.
    A simple mitigation is to provide leaderboards excluding models with potential contamination, and, in addition, a leaderboard that includes all models.
  2. 2.
    Another solution is a healthy dose of suspension by the maintainers. We will generally investigate reasons for better (or worse) performance per dataset, given model outliers. Given that model providers can train on TabArena’s datasets, we expect that for LLM-based approaches, model providers perform memorization tests [122]. For tabular foundation models, we currently have no way of detecting contamination and do not know if such models can “remember” the data in a significant way, as is the case for text- or vision-based models. Here, future research and tools, akin to the work by Bordt et al. [122], are needed.
  3. 3.
    One significant difference between fighting foul play and data contamination in LLM benchmarks, such as ChatbotArena [39], to TabArena is that our official leaderboard requires an open mode (code, data, and result artifacts), which makes potential abuse easier to spot by the community. We do not want to rule out benchmarking API-based closed-source models in the future, so this might also not be a silver bullet.
  4. 4.
    Lastly, we believe that the living benchmark itself will detect foul play across future iterations as new datasets, potentially changing (the seed of) the dataset splits, or new tools to detect foul play will inevitably be added to the TabArena ecosystem.
A shared characteristic among the various future guards is the need to update and maintain the benchmark. Thus, we posit that the strongest protection against foul play is to keep TabArena alive.
Possible Data Contamination in TabArena. It is likely that many benchmarked models were developed or validated on datasets included in TabArena. Since most publications do not disclose this information, the extent of this issue cannot be estimated. Moreover, we do not wish to penalize methods for being transparent about their training and evaluation data. Therefore, we only discuss data contamination and not benchmark overfitting. In TabArena-v0.1, TabDPT is the only model directly pretrained on benchmark datasets, covering seven tasks (KDDCup09_appetency, QSAR-TID-11, Amazon_employee_access, APSFailure, wine_quality, Diabetes130US, heloc). Notably, the model is outperformed by others on most of these datasets, showing a clear advantage only on wine_quality. Although contamination cannot be ruled out entirely, it appears unlikely that the performance of TabDPT is unfairly overestimated compared to others. Therefore, we decided against taking immediate countermeasures. Nevertheless, as contamination could become a serious issue in the future, we will remain cautious when adding models that may have been developed using TabArena datasets.
Benchmarking Foundation Models with Dataset Contamination. Foundation models for tabular data are increasingly trained on real-world data, and often the pretraining data overlaps with the dataset in prior benchmarks; cf. [23, 123, 124, 125]. As a result, we inevitably expect that TabArena will be used to evaluate foundation models that have been trained on some or all of the datasets from TabArena. While we could exclude such models and ignore the problem, this would contradict our goal of benchmarking the state-of-the-art in machine learning on tabular data. Thus, we must consider a future in which we incorporate such models into TabArena. Therefore, we provide a brief discussion on approaches for benchmarking foundation models with dataset contamination below.
The most straightforward approach is to inform practitioners that dataset contamination might exist. For instance, TabArena could follow the lead of GIFT-Eval [56], which recently introduced a boolean “TestData Leakage“ column to the leaderboard. Likewise, one could provide leaderboards with and without models suspected of dataset contamination.
An alternative, more aggressive approach could be to require models that are submitted not to be trained on the datasets from TabArena. The main problem with requiring model developers to provide a “leak-free” version of their foundation models is that we would not benchmark the model used in practice. Thus, our benchmark would not be helpful to practitioners who need to decide between specific models. To explain, consider that we would use a leak-free version as a proxy; then, we need to guarantee that the checkpoint of the leak-free version performs similarly to the original checkpoint, as if one were to change only the pretraining data. However, in deep learning, we generally lack a robust method for pretraining models. In most cases, pretraining requires considerable attention for each training run, such as manually setting learning hyperparameters. Due to the complexity of deep learning, there is no efficient or straightforward way to guarantee that different checkpoints trained on different datasets are representative of each other or achieve maximum performance given their respective datasets. Thus, using a leak-free version as a proxy would most likely not result in accurately benchmarking the best version of the model that practitioners would use.
We conclude that it is debatable how best to benchmark foundation models with dataset contamination. As a minimal measure, TabArena could communicate the presence of dataset contamination. More effective measures require more research. Consequently, we must keep updating TabArena to incorporate more effective measures in the future, assuming future work can identify such measures.
Appendix F Performance Results Per Dataset
This section presents the performance per dataset for all methods in TabArena-v0.1.
@@ Appendix F Performance Results Per Dataset
Table F.1: Performance Per Dataset. We show the average predictive performance per dataset with the standard deviation over folds. We show the performance for the default hyperparameter configuration (Default), for the model after tuning (Tuned), and for the ensemble after tuning (Tuned + Ens.). We highlight the best-performing methods with significance on three levels: (1) Green: The best performing method on average; (2) Bold: Methods that are not significantly worse than the best method on average, based on a Wilcoxon Signed-Rank test for paired samples with Holm-Bonferroni correction and $\alpha=0.05$. (3) Underlined: Methods that are not significantly worse than the best method in the same pipeline regime (Default, Tuned, or Tuned + Ens.), based on a Wilcoxon Signed-Rank test for paired samples with Holm-Bonferroni correction and $\alpha=0.05$. We exclude AutoGluon for significance tests in the Tuned + Ens. regime.
@@
RF 4888.5 $\pm$ 289.2 4641.1 $\pm$ 303.6 4629.6 $\pm$ 302.9
ExtraTrees 4844.9 $\pm$ 274.8 4607.0 $\pm$ 324.0 4609.9 $\pm$ 323.0
XGBoost 4672.0 $\pm$ 306.3 4523.3 $\pm$ 319.7 4519.6 $\pm$ 320.2
LightGBM 4610.4 $\pm$ 313.8 4525.1 $\pm$ 329.2 4511.9 $\pm$ 325.5
CatBoost 4535.4 $\pm$ 328.8 4518.2 $\pm$ 321.8 4519.2 $\pm$ 321.1
EBM 4549.2 $\pm$ 318.8 4499.7 $\pm$ 333.4 4499.3 $\pm$ 325.7
FastAIMLP 4720.2 $\pm$ 313.1 4633.9 $\pm$ 317.7 4624.2 $\pm$ 311.0
TorchMLP 4661.1 $\pm$ 342.8 4526.1 $\pm$ 328.6 4534.0 $\pm$ 333.3
RealMLP 4579.2 $\pm$ 313.2 4570.6 $\pm$ 323.5 4534.8 $\pm$ 323.1
TabM 4513.7 $\pm$ 323.4 4530.0 $\pm$ 323.1 4510.9 $\pm$ 325.7
MNCA 4605.4 $\pm$ 332.0 4614.4 $\pm$ 330.3 4589.2 $\pm$ 331.3
TabPFNv2 4694.7 $\pm$ 302.6 4650.3 $\pm$ 336.9 4567.7 $\pm$ 318.7
TabDPT 4508.4 $\pm$ 295.4 - -